# WiFi-Diffusion: Achieving Fine-Grained WiFi Radio Map Estimation with Ultra-Low Sampling Rate by Diffusion Models

### v1

@@ IV-D Summary
In summary, WiFi-Diffusion is a three-block diffusion model-based algorithm for radio map estimation. The three blocks of WiFi-Diffusion are the boost block, the generation block, and the election block. In the boost block, WiFi-Diffusion utilizes prior information, e.g., layout of buildings, principle of radio propagation, and collected radio samples, to enhance the qualities of radio map candidates generated by the diffusion model. In the generation block, WiFi-Diffusion leverages the creative power of diffusion-based generative model DDIM to produce a diverse set of radio map candidates. In the election block, WiFi-Diffusion employs the mathematical radio propagation model as a guide to find the best map from the diverse candidates generated by the diffusion model. In the following section, we use extensive simulations to verify that WiFi-Diffusion is able to generate high-quality, fine-grained radio maps at ultra-low sampling rates, and show that it significantly outperforms SOTA.
@@ V Evaluation
In this section, we evaluate WiFi-Diffusion. We first use a case study to evaluate its performance and demonstrate its capability of constructing fine-grained, high-quality radio maps at an ultra-low sampling rate. Then we use extensive simulations to compare it against various baseline algorithms, highlighting that it significantly outperforms SOTA at ultra-low sampling rates. Next, we conduct an ablation study to analyze the individual contributions of the boost block and the election block to its overall performance. Finally, we perform a robustness study to evaluate its performance under multiple extreme scenarios for radio map estimation, such as varying numbers of transmitters within the map, transmitters located outside the map, and extremely limited number of radio samples (as few as 3 samples).
@@ V-A Simulation Settings
We use the open-source dataset BART-Lab radio maps [^1] https://github.com/BRATLab-UCD/Radiomap-Data from [28, 29] for evaluation. This dataset is obtained by the simulation software Altair Feko [30] from WinProp and consists of 2000 radio maps. The number of transmitters in one map varies from $1$ to $3$. There are 2000 different layouts of buildings in the dataset, extracted from OpenStreetMap [31] in the United States. The heights of buildings are set as $10$ m. This dataset consists of radio maps simulated in 5 different frequency bands: 1750 MHz, 2750 MHz, 3750 MHz, 4750 MHz, and 5750 MHz. For our evaluation, we specifically use the 5750 MHz (WiFi 5G) data. Given the varying resolutions of the original maps in the dataset, we ensure consistency by randomly selecting regions with a resolution of $256\times 256$ (i.e., $H=W=256$ in simulations) from each map. In total, we have 1776 different radio maps each of which has a resolution of $256\times 256$. For WiFi-Diffusion, we use 1300 maps for training, 376 maps for test, and 100 maps for inference (evaluation), respectively.
@@
We implement WiFi-Diffusion on an Nvidia RTX 4090 GPU and an AMD EPYC 9654 CPU. In WiFi-Diffusion, codes of deep learning are built using the pytorch framework, number of inference steps of DDPM is 1000 ($t=1000$), and number of inference steps of DDIM is 10 ($u=10$). There are 64 map candidates generated by DDIM in the generation block of WiFi-Diffusion for each problem instance ($m=64$).
@@
  • Splines [11]: It uses piecewise low-degree polynomials to approximate the values of unknown points in the curve or surface between known data points.
  • Ordinary kriging [12]: It uses semi-variogram function to describe the spatial autocorrelation and estimates the values of unknown points by weighted average of the observed data of known points.
@@
  • RadioUnet [13]: It employs U-Net for estimating radio maps. U-Net is a convolutional neural network (CNN) architecture, characterized by its symmetric U-shaped structure, consisting of an “encoder” (contracting path) and a “decoder” (expansive path).
  • SkipResidualAutoencoder [15]: It leverages autoencoder for constructing radio maps. Autoencoder is a type of neural network model which is primarily designed to learn effective representations of data. The number of hidden layers in this baseline is set as 128.
  • ResNet [16]: ResNet34 is also use for evaluation. We change the last FC layer to two MLP layers linked with SiLU activation function and activate the output with a sigmoid function. To convert the extracted features to a $256\times 256$ map, the size of the last layer is set as 65536.
To quantify the performance of algorithms, MSE, Root Mean Square Error (RMSE), Normalized Mean Square Error (NMSE), and Peak Signal to Noise Ratio (PSNR) [32] are used.
  • MSE: It is the average of the squared difference between original true values and predicted values.
    $\text{MSE}=\frac{1}{H\times W}\sum^{H-1}_{h=0}\sum^{W-1}_{w=0}[I(h,w)-\hat{I}(h,w)]^{2},$
    where $I(h,w)$ is the original true RSS value at the point $(h,w)$, and $\hat{I}(h,w)$ is the predicted RSS value at that point.
@@
    $\text{RMSE}=\sqrt{\text{MSE}}.$
@@
    $\text{NMSE}=\frac{\displaystyle\sum^{H-1}_{h=0}\sum^{W-1}_{w=0}[I(h,w)-\hat{I}(h,w)]^{2}}{\displaystyle\sum^{H-1}_{h=0}\sum^{W-1}_{w=0}I(h,w)^{2}}.$
@@
    $\text{PSNR}=20\cdot\log_{10}\left(\frac{\text{MAX}_{I}}{\text{RMSE}}\right),$
    where $\text{MAX}_{I}$ is the maximum RSS value of one point in the radio map.
From the above definitions, it is clear that a small MSE, a small RMSE, a small NMSE, or a large PSNR implies that the constructed radio map is of high-quality.
@@ V-B Simulation Results
Case Study    In the case study, we have $k=50$ sensors randomly distributed in one map and collecting radio samples (sampling rate is $0.08\%$). Fig. 10 shows the radio maps constructed by different algorithms, respectively. From the figure, it is clear that the radio map constructed by WiFi-Diffusion is very close to the ground truth (true radio map). It is also clear that the radio map constructed by WiFi-Diffusion is much more accurate than that constructed by any of the 6 baseline algorithms.
More Simulations    Then we perform more simulations to evaluate WiFi-Diffusion. Table III presents the simulation results comparing WiFi-Diffusion with baseline algorithms under varying $k$ that is the number of samples. For each $k$, $100$ radio maps are used for evaluation, and the reported metrics (MSE, RMSE, NMSE, and PSNR) represent the average values obtained from these 100 maps. From the table, we observe that WiFi-Diffusion outperforms all baseline algorithms for every tested value of $k$, spanning from 10 to 200. Moreover, the radio maps generated by WiFi-Diffusion exhibit significantly superior quality compared to those produced by the baseline approaches. Specifically, at ultra-low sampling rates, the MSE values for WiFi-Diffusion can be up to $10$ times lower than (i.e., at most one-tenth of) those of the baseline methods. For example,
@@
  • WiFi-Diffusion vs Ordinary kriging: The MSE of WiFi-Diffusion is $0.0008$, which is over 15 times lower than $0.0126$ that is the MSE of ordinary kriging when $k=100$.
@@
Fig. 11 presents curves of the simulation results comparing WiFi-Diffusion with the three deep learning-based baseline algorithms under varying $k$. From the figure, we observe that the MSE achieved by WiFi-Diffusion converges to a value close to 00 when $k=50$. This implies that WiFi-Diffusion generates fine-grained, high-quality radio maps at an ultra-low sampling rate of $50/(256\times 256)=0.08\%$. In comparison, the converged MSE of ResNet is significantly greater than that of WiFi-Diffusion. For RadioUNet and SkipResidualAutoencoder, they converge when $k=150$, resulting in a convergence speed much slower than that of WiFi-Diffusion. Similar conclusions can also be derived for RMSE, NMSE, and PSNR, respectively, based on the results in Fig. 11.
@@
Ablation Study    WiFi-Diffusion is a three-block algorithm. In addition to the important diffusion model-based generation block, it has a boost block to optimize the performance of the generation block and an election block to find the best map from the set of maps produced by the generation block. Now we use an ablation study to evaluate the boost block and the election block, respectively, of WiFi-Diffusion.
Table IV gives the simulation results of the ablation study, which clearly demonstrate the necessity of both the boost block and the election block in WiFi-Diffusion. The performance of WiFi-Diffusion is significantly degraded when either the boost block or the election block is removed, highlighting their critical roles in the overall functionality of WiFi-Diffusion.
TABLE IV: Ablation study to evaluate the boost block and election block of WiFi-Diffusion, respectively.
Setting MSE under different number of samples
@@
Fig. 12 gives radio maps generated by WiFi-Diffusion and three deep learning-based baseline algorithms, under scenarios with differing numbers of transmitters. Specifically, Fig. LABEL:r-1-1 presents results with a single transmitter, while Fig. LABEL:r-1-2 illustrates results with two transmitters, both with $k=50$. The results demonstrate that WiFi-Diffusion consistently produces high-quality radio maps, regardless of the number of transmitters, and significantly outperforms the baseline algorithms.
@@
Furthermore, Fig. 13 evaluates the robustness of WiFi-Diffusion when transmitters are positioned outside the map. In Fig. LABEL:r-2-1, no transmitters are located within the map, but two are positioned outside, whereas in Fig. LABEL:r-2-2, two transmitters are within the map and one outside, again with $k=50$. Fig. 13 indicates that WiFi-Diffusion is able to construct high-quality radio maps even when certain transmitters are located outside the map, significantly outperforming baseline algorithms.
@@
Additionally, we evaluate WiFi-Diffusion in an extremely challenging scenario where the sampling rate is nearly 00, specifically when $k=3$. The results of this evaluation are presented in Fig. 14. Despite the daunting task of predicting RSS values for a $256\times 256$ grid using only $3$ known samples, WiFi-Diffusion manages to produce a rough radio map. In contrast, all six baseline algorithms fail to generate even a basic representation of the radio map. This comparison underscores the exceptional robustness of WiFi-Diffusion in estimating radio maps, particularly in extremely challenging scenarios when the sampling rate is nearly 00.
@@
Figure 14: Radio maps constructed by different algorithms when there are only $3$ radio samples.
VI Conclusion

### v3

@@ IV-D Summary
In summary, WiFi-Diffusion is a three-block diffusion model-based algorithm for fine-grained radio map estimation. The three blocks of WiFi-Diffusion are the boost block, the generation block, and the election block. In the boost block, WiFi-Diffusion utilizes prior information, e.g., layout of buildings, principles of radio propagation, and collected radio samples, to enhance the qualities of fine-grained radio map candidates generated by the diffusion model. In the generation block, WiFi-Diffusion leverages the creative power of the diffusion-based generative model to produce a diverse set of fine-grained radio map candidates. In the election block, WiFi-Diffusion uses the mathematical radio propagation model as a guide to find the best fine-grained map from the diverse candidates generated by the diffusion model. In the following section, we use extensive simulations to verify that WiFi-Diffusion is able to generate fine-grained, high-quality radio maps at ultra-low sampling rates, and show that it significantly outperforms SOTA.
@@ V Evaluation
In this section, we evaluate WiFi-Diffusion. We first use a case study to evaluate its performance and demonstrate its capability of constructing fine-grained, high-quality radio maps at ultra-low sampling rates. Then we use extensive simulations to compare it against various baseline algorithms, highlighting that it significantly outperforms SOTA at ultra-low sampling rates. Next, we conduct ablation studies to analyze the individual contributions of the boost block and the election block to its overall performance. In addition, we perform robustness studies to evaluate its performance under multiple extreme scenarios for radio map estimation, such as varying numbers of transmitters on one map, transmitters located outside the map, and extremely limited number of collected radio samples (as few as 3 samples). Finally, we present its model size and training cost, and use simulations to show that it maintains significant performance advantages over SOTA, even when SOTA is scaled up to comparable or greater complexity and computational expense.
@@ V-A Simulation Settings
We use the open-source dataset BART-Lab radio maps [^1] https://github.com/BRATLab-UCD/Radiomap-Data from [27, 28] for the evaluation. This dataset is obtained by the simulation software Altair Feko [29] from WinProp and consists of 2000 radio maps. The number of transmitters on one map varies from $1$ to $3$. There are 2000 different layouts of buildings in the dataset, extracted from OpenStreetMap [30] in the United States. The heights of the buildings are set as $10$ m. This dataset consists of radio maps simulated in 5 different frequency bands: 1750 MHz, 2750 MHz, 3750 MHz, 4750 MHz, and 5750 MHz. For our evaluation, we specifically use the 5750 MHz (WiFi 5G) data. Considering that the resolutions of the original maps in the dataset are different, to ensure consistency, we randomly select regions with a resolution of $256\times 256$ (i.e., $H=W=256$) from each map. In total, we have 1776 different radio maps, each of which has a resolution of $256\times 256$. For WiFi-Diffusion, we use 1300 maps for training, 376 maps for testing, and 100 maps for inference (evaluation), respectively.
@@
We implement WiFi-Diffusion on an NVIDIA RTX 4090 GPU and an AMD EPYC 9654 CPU. For WiFi-Diffusion, deep learning codes are built using PyTorch, the number of DDPM inference steps is 1000 ($t=1000$), and the number of DDIM inference steps is 10 ($u=10$). There are 64 map candidates generated by DDIM in the WiFi-Diffusion generation block for each problem instance ($m=64$).
@@
  • Splines [11]: It uses piecewise low-degree polynomials to approximate the values of unknown points on the curve or surface between known data points.
  • Ordinary kriging [12]: It uses a semi-variogram function to describe the spatial autocorrelation and estimates the values of unknown points by weighted average of the observed data of known points.
@@
  • RadioUnet [13]: It uses UNet to estimate radio maps. UNet is a convolutional neural network (CNN) architecture, characterized by its symmetric U-shaped structure, consisting of an “encoder” (contracting path) and a “decoder” (expansive path).
  • SkipResidualAutoencoder [15]: It uses an auto-encoder to build radio maps. The autoencoder is a type of neural network model that is primarily designed to learn effective representations of data. The number of hidden layers in this baseline is set to 128.
  • ResNet [16]: ResNet34 is also used for evaluation. We change the last FC layer to two MLP layers linked with the SiLU activation function and activate the output with a sigmoid function. To convert the extracted features to a $256\times 256$ map, the size of the last layer is set to 65536.
To quantify the performance of algorithms, MSE, Root Mean Square Error (RMSE), Normalized Mean Square Error (NMSE), and Peak Signal to Noise Ratio (PSNR) [31] are used:
  • MSE: It is the average of the squared difference between the original true values and the predicted values.
    $\text{MSE}=\frac{1}{H\times W}\sum^{H-1}_{h=0}\sum^{W-1}_{w=0}[I(h,w)-\hat{I}(h,w)]^{2},$ (27)
    where $I(h,w)$ is the original true RSS value at the point $(h,w)$ and $\hat{I}(h,w)$ is the predicted RSS value at that point.
@@
    $\text{RMSE}=\sqrt{\text{MSE}}.$ (28)
@@
    $\text{NMSE}=\frac{\displaystyle\sum^{H-1}_{h=0}\sum^{W-1}_{w=0}[I(h,w)-\hat{I}(h,w)]^{2}}{\displaystyle\sum^{H-1}_{h=0}\sum^{W-1}_{w=0}I(h,w)^{2}}.$ (29)
@@
    $\text{PSNR}=20\cdot\log_{10}\left(\frac{\text{MAX}_{I}}{\text{RMSE}}\right),$ (30)
    where $\text{MAX}_{I}$ is the maximum RSS value of a point on the radio map.
From the above definitions, it is clear that a small MSE, a small RMSE, a small NMSE, or a large PSNR implies that the constructed radio map is of high quality.
@@ V-B Simulation Results
Case Study    In the case study, we have $k=50$ sensors randomly distributed on a map (the sampling rate is $0.08\%$). Fig. 10 shows the radio maps constructed by different algorithms, respectively. From the figure, it is clear that the radio map constructed by WiFi-Diffusion is very close to the ground truth (true radio map). It is also clear that the radio map constructed by WiFi-Diffusion is much more accurate than that constructed by any of the 6 baseline algorithms.
More Simulations    Then we perform more simulations to evaluate WiFi-Diffusion. Table III presents the simulation results comparing WiFi-Diffusion with the baseline algorithms under varying $k$ (i.e., number of samples). For each $k$, $100$ radio maps are used for evaluation, and the reported metrics (MSE, RMSE, NMSE, and PSNR) represent the average values obtained from these 100 maps. From the table, we observe that WiFi-Diffusion outperforms all baseline algorithms for every value of $k$, ranging from 10 to 200. Moreover, the radio maps generated by WiFi-Diffusion exhibit significantly superior quality compared to those produced by the baseline approaches. Specifically, at ultra-low sampling rates, the MSE values for WiFi-Diffusion can be up to $10$ times lower than (i.e., at most one-tenth of) those of the baseline methods. For example,
@@
  • WiFi-Diffusion vs Ordinary kriging: The MSE of WiFi-Diffusion is $0.0008$, which is over 15 times lower than $0.0126$ which is the MSE of ordinary kriging when $k=100$.
@@
Fig. 11 presents curves of simulation results that compare WiFi-Diffusion with the three deep learning-based baseline algorithms under varying $k$. From the figure, we observe that the MSE achieved by WiFi-Diffusion is close to 00 when $k\geq 50$. This implies that WiFi-Diffusion generates fine-grained, high-quality radio maps at sampling rates lower than $0.1\%$. In comparison, the MSE of ResNet is significantly greater than 00 for any $k$. The MSE values of RadioUnet and SkipResidualAutoencoder are close to 00 when $k\geq 150$, but are much greater than 00 when $k\leq 100$. As a result, none of the three baseline algorithms can generate fine-grained, high-quality radio maps at sampling rates lower than $0.1\%$. Similar conclusions can also be drawn for RMSE, NMSE, and PSNR, respectively, based on the results in Fig. 11.
Ablation Study    WiFi-Diffusion consists of three blocks. In addition to the important diffusion model-based generation block, it has a boost block to optimize the performance of the generation block and an election block to find the best map from the set of maps produced by the generation block. Now we use ablation studies to evaluate the boost block and the election block, respectively, of WiFi-Diffusion.
TABLE IV: Ablation studies to evaluate different blocks of WiFi-Diffusion.
Setting MSE under varying number of samples
@@
TABLE V: Simulation results to verify the effectiveness of the election block of WiFi-Diffusion.
Setting MSE under varying number of samples
10 25 50
No election block $0.01246$ $0.00392$ $0.00199$
With an election block that minimizes (31) $0.00970$ $0.00292$ $0.00138$
With our proposed election block defined by (26) $\boldsymbol{0.00796}$ $\boldsymbol{0.00277}$ $\boldsymbol{0.00133}$
100 150 200
No election block $0.00139$ $0.00125$ $0.00128$
With an election block that minimizes (31) $0.00093$ $0.00073$ $0.00068$
With our proposed election block defined by (26) $\boldsymbol{0.00082}$ $\boldsymbol{0.00070}$ $\boldsymbol{0.00063}$
Table IV gives the simulation results of the ablation studies, which clearly demonstrate the necessity of the boost block and the election block in WiFi-Diffusion. The performance of WiFi-Diffusion is significantly degraded when either the boost block or the election block is removed, highlighting their critical roles in the overall functionality of WiFi-Diffusion.
Figure 12: The set of $64$ radio map candidates (as well as the achieved MSE values) generated by the generation block of WiFi-Diffusion in one simulation instance when $k=25$ (the red map is the map identified by the election block and is the best map in the set; the blue map is the map with the lowest $\text{MSE}_{r}$ defined in (31) and is much worse than the red map).
In the following, we use simulations to further verify the effectiveness of the election block. Considering that the true radio map (map to be predicted) is unknown during inference, different from our proposed design of the election block, one straightforward design is to identify the candidate map with predicted RSS values closest to the radio samples collected by the sensors. Specifically, we can calculate $\text{MSE}_{r}$ defined as follows:
$\text{MSE}_{r}~{}=~{}\frac{1}{k}\sum^{k}_{i=1}[I_{i}-\hat{I}_{i}]^{2},$ (31)
where $I_{i}$ represents the measured RSS value from sensor $i$ and $\hat{I}_{i}$ denotes the predicted RSS value at the corresponding location in the generated map. The optimal map is identified as the candidate that minimizes $\text{MSE}_{r}$.
Fig. 12 illustrates a representative simulation instance, displaying the candidate set of 64 fine-grained radio maps generated by the generation block of WiFi-Diffusion along with their corresponding MSE values. It demonstrates a significant variation in map quality, with the best map achieving an MSE of $0.0013$ compared to $0.0040$ for the poorest map, indicating that it is necessary to have an election block in WiFi-Diffusion. Our proposed election design successfully identifies the optimal map (red map) with the lowest MSE of $0.0013$, while the simple election design based solely on the minimization of $\text{MSE}_{r}$ selects an inferior candidate (blue map) with an MSE of $0.0021$.
Moreover, we simulate more instances under varying $k$ that is the number of samples, with the results summarized in Table V. The results consistently show that our election block design outperforms the straightforward $\text{MSE}_{r}$-based baseline, particularly when $k=10$, $k=25$, and $k=100$. In general, Fig. 12 and Table V demonstrate both the necessity of incorporating an election block in WiFi-Diffusion and the superior performance of our proposed design over alternative approaches.
The election block can improve the performance of WiFi-Diffusion by using the law of radio propagation as a guide to select the high-quality radio map from a set of candidate maps produced by the generation block. The three-block mechanism of WiFi-Diffusion stems from the inherent strength of generative AI approaches like diffusion models in producing diverse, innovative solutions. WiFi-Diffusion effectively leverages this generative advantage, which is in contrast to conventional discriminative models such as RadioUnet, SkipResidualAutoencoder, and ResNet. Although these discriminative approaches offer stable predictions, their nature fundamentally limits their ability to generate diverse output variations. Consequently, it is difficult to add similar election blocks to them to improve their performance.
@@
Fig. 13 shows radio maps generated by WiFi-Diffusion and three deep learning-based baseline algorithms, under scenarios with different numbers of transmitters. Specifically, Fig. LABEL:r-1-1 presents results with a single transmitter, while Fig. LABEL:r-1-2 illustrates results with two transmitters, both with $k=50$. The results demonstrate that WiFi-Diffusion consistently produces fine-grained, high-quality radio maps, regardless of the number of transmitters, and significantly outperforms the baseline algorithms.
@@
Furthermore, Fig. 14 evaluates the robustness of WiFi-Diffusion when certain transmitters are located outside the map. In Fig. LABEL:r-2-1, no transmitters are located within the map but two outside, while in Fig. LABEL:r-2-2, two transmitters are located within the map and one outside, again with $k=50$. Fig. 14 indicates that WiFi-Diffusion is able to construct fine-grained, high-quality radio maps even when certain transmitters are located outside the map, significantly outperforming baseline algorithms.
@@
Additionally, we evaluate WiFi-Diffusion in an extremely challenging scenario where the sampling rate is nearly 00, specifically when $k=3$. The results of this evaluation are presented in Fig. 15. Despite the daunting task of predicting RSS values for a $256\times 256$ grid using only $3$ known samples, WiFi-Diffusion manages to produce a rough radio map. In contrast, all six baseline algorithms fail to generate even a basic representation of the radio map. This comparison underscores the exceptional robustness of WiFi-Diffusion in estimating radio maps, particularly in the extremely challenging scenario when the sampling rate is nearly 00.
@@
Figure 15: Radio maps constructed by different algorithms when there are only $3$ collected radio samples.
TABLE VI: Number of parameters, time of inference, and time of training of different algorithms (m: million; s: second; h: hour).
RadioUnet Autoencoder ResNet WiFi-Diffusion
Number of
parameters
13m 4m 89m 122m
Time of
inference
0.06s 0.06s 0.07s 7.82s
Time of
training
19h 7h 11h 44h
Model Size and Training Cost    Due to their inherent complexity, diffusion models typically require more parameters and longer computational times compared to traditional deep learning architectures such as UNet. As shown in Table VI, which compares the model size (number of parameters)222 In Table VI, the number of parameters of WiFi-Diffusion is near $122$ million, where the number of parameters for the boost block is near $61$ million and that for the generation block is near $61$ million., inference time, and training time across different algorithms, WiFi-Diffusion is much larger and slower than the baseline methods. A critical question arises:
Would baseline methods perform comparably or even outperform WiFi-Diffusion if their parameter count were increased to match or exceed that of WiFi-Diffusion?
We address this question empirically through comprehensive evaluations.
We enhance each of the three baseline models – RadioUnet, Autoencoder, and ResNet – by increasing their depth (number of layers), expanding their width (channel dimensions), and training them with parameter counts and computational time comparable to or exceeding those of WiFi-Diffusion. The resulting improved variants are denoted as RadioUnet+, Autoencoder+, and ResNet+, respectively. As demonstrated in Table VII, which summarizes the model size (number of parameters), inference time, and training time across all evaluated algorithms, the enhanced baselines (RadioUnet+, Autoencoder+, and ResNet+) exhibit larger model sizes and longer training durations than WiFi-Diffusion.
Figure 16: Simulation results of MSE and inference time achieved by different algorithms.
Fig. 16 compares the inference times across different algorithms and presents the ratio between the MSE values achieved by WiFi-Diffusion and those achieved by the baseline methods. Table VII also shows the results of inference times and MSE values achieved by different algorithms. From Fig. 16 and Table VII we observe that (i) the radio map constructed by WiFi-Diffusion is significantly better than that constructed by any of the enhanced baselines (RadioUnet+, Autoencoder+, and ResNet+), considering that the MSE achieved by WiFi-Diffusion is $70\%$ lower than that achieved by any of the enhanced baselines when $k=25$ and the MSE achieved by WiFi-Diffusion is $60\%$ lower than that achieved by any of the enhanced baselines when $k=50$; (ii) While WiFi-Diffusion requires a longer inference time than any of the enhanced baselines, the absolute inference time of WiFi-Diffusion remains practical and acceptable, never exceeding $10$ seconds, enabling real-time radio map generation. Moreover, it is worth highlighting that manually increasing the inference time for any of the enhanced baseline methods does not yield improvements in the quality of constructed radio maps.
TABLE VII: Number of parameters, time of inference, time of training, and MSE results of different algorithms (m: million; s: second; h: hour).
Method Number of parameters Inference time MSE ($k=25$) MSE ($k=50$) Training time
RadioUnet 13,274,031 (13m) 0.06s 0.0450 0.0159 19h
RadioUnet+ 144,249,256 (144m) 0.44s 0.0153 0.0032 48h
Autoencoder 3,553,945 (4m) 0.06s 0.0239 0.0071 7h
Autoencoder+ 127,491,225 (127m) 0.11s 0.0123 0.0042 55h
ResNet 88,981,248 (89m) 0.07s 0.0104 0.0092 11h
ResNet+ 161,984,512 (162m) 0.11s 0.0101 0.0091 45h
WiFi-Diffusion 60,823,169 $\times$ 2 (122m) 7.82s 0.0027 0.0013 44h
V-C Summary
In summary, the evaluations given in this section demonstrate that:
  • WiFi-Diffusion is capable of generating fine-grained, high-quality radio maps when the sampling rate is less than $0.1\%$.
  • WiFi-Diffusion achieves superior accuracy, yielding radio maps with MSE values an order of magnitude lower than those produced by the baseline methods under same ultra-low sampling rates.
  • WiFi-Diffusion achieves comparable accuracy (MSE) to the baseline methods while requiring only $20\%$ of their sampling rate for fine-grained radio map generation.
  • All three blocks including the boost block, generation block, and election block are essential for WiFi-Diffusion’s high-performance on radio map estimation.
  • Even when the baseline methods are scaled up to match or exceed WiFi-Diffusion’s model size and training cost, WiFi-Diffusion maintains significantly superior accuracy, achieving radio maps with MSE values over $60\%$ lower than those produced by the baseline methods.
VI WiFi-Diffusion: Potential Applications
WiFi-Diffusion generates fine-grained, high-quality radio maps which present parameters of interest in communication channels, such as the RSS, at every point of a certain geographical region for the WiFi spectrum. It can improve the performance of many WiFi-based wireless networks, including unmanned aerial vehicle (UAV) network, industrial Internet of Things (IoT), and Internet of Vehicles (IoV).
UAV Network    It serves as the critical communication infrastructure which enables wireless connectivity and information exchange among multiple drones [32]. Modern UAVs equipped with onboard sensors are capable of collecting and transmitting radio samples. An edge server can leverage these collected samples through WiFi-Diffusion to construct and disseminate fine-grained, accurate radio maps to individual UAVs in real-time. These radio maps provide UAVs with predictive capabilities for RSS values in various frequency bands along their designated flight trajectories. Consequently, each UAV can dynamically optimize its spectrum access strategy by adaptively adjusting transmission parameters according to spatial and temporal variations. In this case, WiFi-Diffusion can be used to improve key performance metrics, such as throughput, for each UAV’s data transmission across its flight path.
Industrial IoT    It is the network of interconnected smart devices, sensors, and machinery in industrial environments that collect, analyze, and act on data to optimize manufacturing, logistics, and other industrial operations [33]. WiFi-Diffusion enables the generation of fine-grained, high-quality radio maps, providing spectrum utilization analytics across spatial, temporal, and frequency domains within the industrial environment. These insights facilitate dynamic optimization of the placement and spectrum access strategies of smart devices, sensors, and machinery. By implementing WiFi-Diffusion, critical industrial communication metrics, particularly latency and reliability, can be systematically enhanced, ensuring timely and robust data transmission for mission-critical industrial operations.
IoV    It is an intelligent ecosystem that connects vehicles, infrastructure, and users through Vehicle-to-Everything (V2X) communication, enabling real-time safety alerts, traffic optimization, and autonomous driving [34]. The reliance on open wireless protocols makes the system susceptible to intentional jamming attacks and unintentional interference. To mitigate these vulnerabilities, fine-grained, high-quality radio maps generated by WiFi-Diffusion can provide a critical defense mechanism by delivering dynamic spectrum awareness and enabling prediction of interference patterns across temporal, spatial, and frequency domains. These radio maps endow IoV with anti-jamming capabilities and enable IoV to adaptively reconfigure communication strategies, e.g., proactive channel switching to uncontaminated frequencies.
VII Conclusion